OpenStudy (vishweshshrimali5):

CHALLENGE QUESTION 5 The vertices of a triangle are (1,0,0),(0,\(\cfrac{1}{\sqrt{2}}\), 0), (0,0,\(\cfrac{1}{\sqrt{3}}\)). If its orthocentre is (\(\alpha,\beta,\gamma\)), then \(\large{(\alpha\beta\gamma)^{\huge{\frac{-2}{5}}}}\) is the answer ranges from 0 to 9...

OpenStudy (vishweshshrimali5):

@waterineyes

OpenStudy (vishweshshrimali5):

The vertices of a triangle are (1,0,0),(0,\(\cfrac{1}{\sqrt{2}}\), 0), (0,0,\(\cfrac{1}{\sqrt{3}}\)). If its orthocentre is (\(\alpha,\beta,\gamma\)), then \(\large{(\alpha\beta\gamma)^{\huge{\frac{-2}{5}}}}\) is the answer ranges from 0 to 9... This is the question

OpenStudy (anonymous):

It should be divided by 3..

OpenStudy (anonymous):

I think I am getting: \[\color{green}{3 \times \sqrt[5]{18}}\]

OpenStudy (vishweshshrimali5):

How did u get that @completeidiot

OpenStudy (vishweshshrimali5):

\(\huge\color{green}{3 \times\sqrt[5]{18}}\)

OpenStudy (vishweshshrimali5):

This is more clearly visible !!

OpenStudy (vishweshshrimali5):

Random guess or any trick...... ?

OpenStudy (anonymous):

Orthocenter is given by Average of the respective coordinates.

OpenStudy (vishweshshrimali5):

no........... @waterineyes

OpenStudy (anonymous):

my method was incorrect, ignore it

OpenStudy (vishweshshrimali5):

that is what we call centroid @waterineyes

OpenStudy (vishweshshrimali5):

All right @completeidiot

OpenStudy (anonymous):

Oh Then messed up sorry..

OpenStudy (vishweshshrimali5):

np

OpenStudy (anonymous):

Here we have to do more..

OpenStudy (vishweshshrimali5):

that's why it is a challenge problem :)

OpenStudy (anonymous):